OpenStudy (anonymous):

lim x->0 [(3sin4x)/(sin 3x)]

OpenStudy (tkhunny):

Super big hint: \(\dfrac{3\sin(4x)}{\sin(3x)} = \dfrac{3x\sin(4x)}{x\sin(3x)} = 4\cdot\dfrac{3x\sin(4x)}{4x\sin(3x)} = 4\cdot\dfrac{\sin(4x)}{4x}\cdot\dfrac{3x}{\sin(3x)}\)